OpenStudy (anonymous):

Find the length of the curve y=x^(2/3) x=0, x=3

OpenStudy (anonymous):

\[\int\limits_{0}^{3}\sqrt{1+(4/9)x^{-2/3}}dx\]

OpenStudy (freckles):

write in terms of y instead your adventure should be much easier

OpenStudy (freckles):

\[x=y^\frac{3}{2} , y=0 \text{ \to } y=3^\frac{2}{3} \\ x'=\frac{3}{2}y^\frac{1}{2} \\ (x')^2=\frac{9}{4}y \\ \int\limits_0^{3^{\frac{2}{3}}} \sqrt{1+\frac{9}{4} y} dy\] use a substitution for the inside v=1+9y/4

OpenStudy (anonymous):

thanks you were right, it made things so much easier

OpenStudy (freckles):

http://tutorial.math.lamar.edu/Classes/CalcII/ArcLength.aspx Pauls notes does a similar one if you wanted more examples. And I'm glad it helped you. :)